# ChartM3: Benchmarking Chart Editing with Multimodal Instructions

### v1

@@ Abstract.
Charts are a fundamental visualization format widely used in data analysis across research and industry. While enabling users to edit charts based on high-level intentions is of great practical value, existing methods primarily rely on natural language instructions, which are often too ambiguous to support fine-grained editing. In this work, we introduce a novel paradigm for multimodal chart editing, where user intent is expressed through a combination of natural language and visual indicators that explicitly highlight the elements to be modified. To support this paradigm, we present Chart, a new benchmark for Multimodal chart editing with Multi-level complexity and Multi-perspective evaluation. Chart contains 1,000 samples spanning four levels of editing difficulty. Each sample includes triplets in the form of (chart, code, multimodal instructions). To comprehensively evaluate chart editing models, Chart provides metrics that assess both visual appearance and code correctness. Our benchmark reveals significant limitations in current multimodal large language models (MLLMs), including GPT-4o, particularly in their ability to interpret and act on visual indicators. To address this, we construct Chart-Train, a large-scale training set with 24,000 multimodal chart editing samples. Fine-tuning MLLMs on this dataset leads to substantial improvements, demonstrating the importance of multimodal supervision in building practical chart editing systems. Our datasets, codes, and evaluation tools are available at https://github.com/MLrollIT/ChartM3.
@@ 2. Related Works
2.1. Chart-related Research.
Charts, as a crucial form of data visualization, have garnered significant research interest, with efforts focusing on both chart understanding and generation. Chart understanding tasks such as chart question answering (Chen et al., 2025; Masry et al., 2022; Methani et al., 2019; Kantharaj et al., 2022; Vogel et al., 2025), chart-to-table conversion, and chart captioning aim to comprehend the semantics of chart content. In contrast, chart generation tasks like text-to-chart (Pesaran Zadeh et al., 2024) and chart-to-code (Zhao et al., 2025) focus on automating chart creation. Recent studies (Yan et al., 2024; Goswami et al., 2025; Yang et al., 2025) introduced chart editing, which allows for interactive chart modification. However, these works primarily rely on natural language instructions, which often struggle to represent fine-grained user intentions. Our proposed Chart introduces visual indicators for more precise chart element manipulation and evaluates a model’s understanding of both chart images and their code.
2.2. Image Editing
Image editing has evolved from pixel-level tasks like inpainting (Pathak et al., 2016) and style transfer (Ghiasi et al., 2017) to text-based (Kawar et al., 2022; Sheynin et al., 2023) and drag-based (Pan et al., 2023; Shi et al., 2023) techniques using diffusion models (Sohl-Dickstein et al., 2015) and GANs (Goodfellow et al., 2014). These methods excel in modifying natural and artistic images. However, chart editing is more complex, requiring models to preserve unedited elements while transforming specific visual components based on code. This makes chart editing a unique challenge, demanding models to translate visual intentions into functional code.
2.3. Multimodal Large Language Models.
Multimodal large language models (MLLMs) have made substantial progress, with closed-source models like GPT-4V (OpenAI, 2023) excelling in real-world vision-language tasks. Open-source models like LLaVA (Liu et al., 2023), Qwen-VL (Wang et al., 2024), and DeepSeek-VL (Lu et al., 2024) have demonstrated strong performance in multimodal understanding benchmarks. However, these models often lag behind closed-source alternatives in practical applications. Our ChartM3 benchmark addresses this gap by evaluating MLLMs on chart editing, using both visual indicators and textual instructions, providing a more user-centric evaluation framework that aligns with real-world use cases.
3. The ChartM3 Benchmark
@@ 3.4. Multi-perspective Evaluation Metrics
SSIM directly quantifies the visual‐quality improvements introduced by the editing process by measuring the proportional gain in structural similarity between the generated output and the target, relative to the initial input’s similarity. In contrast, the GPT Score operates at the semantic level—evaluating directive compliance and non-intervened robustness—to capture the model’s reasoning and code-generation capabilities. Together, these two metrics offer complementary insights, with SSIM assessing perceptual fidelity and GPT Score probing the correctness and precision of the underlying modification logic.
@@ 3.5. Supervised Fine-Tuning
Since existing chart editing (Chen et al., 2025; Yang et al., 2025; Masry et al., 2022; Yan et al., 2024; Zhao et al., 2025; Pesaran Zadeh et al., 2024; Vogel et al., 2025) frameworks typically adopts the traditional text-guided editing paradigm, they can hardly support multimodal chart editing used in ChartM3. To improve the multimodal chart editing capabilities of the model, we construct ChartM3-Train dataset that consists of 24,000 samples with the same data construction pipeline described in Section 3.2. ChartM3-Train also contains same amount of visual indicator-guided samples and textual descrption-guided samples.
The model processes instructions (either basic or descriptive), images (annotated or unannotated), and initial code as input, and is trained to generate reference code as output. We used cross-entropy loss to minimize discrepancies between the generated code and the reference code. This approach enables our model to simultaneously develop two capabilities: interpreting visual annotations for precise code modifications and understanding textual references to chart elements without explicit visual guidance.

### v2

@@ Abstract.
Charts are a fundamental visualization format widely used in data analysis across research and industry. While enabling users to edit charts based on high-level intentions is of great practical value, existing methods primarily rely on natural language instructions, which are often too ambiguous to support fine-grained editing. In this work, we introduce a novel paradigm for multimodal chart editing, where user intent is expressed through a combination of natural language and visual indicators that explicitly highlight the elements to be modified. To support this paradigm, we present Chart, a new benchmark for Multimodal chart editing with Multi-level complexity and Multi-perspective evaluation. Chart contains 1,000 samples spanning four levels of editing difficulty. Each sample includes triplets in the form of (chart, code, multimodal instructions). To comprehensively evaluate chart editing models, Chart provides metrics that assess both visual appearance and code correctness. Our benchmark reveals significant limitations in current multimodal large language models (MLLMs), including GPT-4o, particularly in their ability to interpret and act on visual indicators. To address this, we construct Chart-Train, a large-scale training set with 24,000 multimodal chart editing samples. Fine-tuning MLLMs on this dataset leads to substantial improvements, demonstrating the importance of multimodal supervision in building practical chart editing systems.
@@ 2. Related Works
Chart-related Research. As a specific image specific type of image for data visualization, charts have attracted increasing research attention. Existing works explore both chart understanding and generation tasks. For example, chart understanding tasks such as chart question answering (Chen et al., 2025; Masry et al., 2022; Methani et al., 2019; Kantharaj et al., 2022; Vogel et al., 2025), chart-to-table, and chart captioning mainly focus on semantic understanding of chart content. Chart generation tasks such as text-to-chart (Pesaran Zadeh et al., 2024) and chart-to-code (Zhao et al., 2025) generation focus on automatic chart creation. Recent works (Yan et al., 2024; Goswami et al., 2025; Yang et al., 2025) propose the chart editing task that enables modification of specific chart elements based on user intention. It enables interactive chart creation with users. Though progress has been made, these works represent user intention primarily with natural language instructions, which can hardly identify fine-grained manipulation of specific chart elements due to ambiguity. Apart from natural language instructions, our proposed Chart provides visual indicators to precisely identify the elements to be edited. It also offers a fresh perspective to evaluate a model’s understanding of both chart images and their corresponding code.
Image Editing. Image editing evolves from pixel-level manipulation tasks, such as inpainting (Pathak et al., 2016) and style transfer (Ghiasi et al., 2017), to more advanced text-based (Kawar et al., 2022; Sheynin et al., 2023) and drag-based (Pan et al., 2023; Shi et al., 2023) editing based on diffusion models (Sohl-Dickstein et al., 2015) and generative adversarial networks (GAN) (Goodfellow et al., 2014). These approaches primarily focus on making visual modifications based on textual prompts or user-provided drag instructions. They excel in editing natural images or artistic images, producing creative and consistent results. In contrast to pixel-level editing, chart editing requires the model to strictly preserve visual elements that are not specified to be edited, and thus requires editing based on its underlying code. It thus presents an additional challenge to the model to translate the visual intention into appropriate code modification.
Multimodal Large Language Models. Research in Multimodal large language models (MLLMs) achieved significant progress in recent years. For example, close-sourced MLLMs such as GPT-4V (OpenAI, 2023) is capable of completing complex real world vision and language tasks. Open-source MLLMs such as LLaVA (Liu et al., 2023; Xu et al., 2024), Qwen-VL (Wang et al., 2024), and DeepSeek-VL (Lu et al., 2024) also demonstrate good performance compared to previous method in many standard multimodal understanding benchmarks. However, these models frequently underperform close-sourced models in real-world scenarios. Thus, it is necessary to construct new evaluation frameworks that reflect practical applications. Our ChartM3 benchmark resolves this demand by evaluating MLLMs’ capability with chart editing based on visual indicators and textual instructions. It assesses the capacity to perform visual language understanding and generate functional code, and provides a more user-centric evaluation setting that closely aligns with real-world usage scenarios.
3. The ChartM³ Benchmark
@@ 3.4. Multi-perspective Evaluation Metrics
SSIM directly quantifies the visual-quality improvements introduced by the editing process by measuring the proportional gain in structural similarity between the generated output and the target, relative to the initial input’s similarity. In contrast, the GPT Score operates at the semantic level-evaluating directive compliance and non-intervened robustness-to capture the model’s reasoning and code-generation capabilities. Together, these two metrics offer complementary insights, with SSIM assessing perceptual fidelity and GPT Score probing the correctness and precision of the underlying modification logic.
@@ 3.5. Supervised Fine-Tuning
Existing chart editing frameworks (Chen et al., 2025; Yang et al., 2025; Masry et al., 2022; Yan et al., 2024; Zhao et al., 2025; Pesaran Zadeh et al., 2024; Vogel et al., 2025) rely on text-guided editing, limiting support for multimodal chart editing like in ChartM3. To enhance its multimodal capabilities, we create the ChartM3-Train dataset, which includes 24,000 samples using the same data construction pipeline from Section 3.2. This dataset contains an equal number of visual indicator-guided and textual description-guided samples.
The model processes instructions (basic or descriptive), images (annotated or unannotated), and initial code as input, generating reference code as output. We apply cross-entropy loss to minimize discrepancies between the generated and reference code, enabling the model to interpret visual annotations and textual references for accurate code modifications.
@@ 6. Conclusion
Acknowledgements
We thank all reviewers for their insightful comments and suggestions. This work was partially supported by the Beijing Natural Science Foundation (No. L233008).